OpenStudy (anonymous):

Casey and Jeremiah made a simple bike ramp out of a piece of straight plywood that is thirty inches long. If the top of the ramp stands fourteen inches high, what is the length of the ramp along the ground? Round your answer to the nearest inch

OpenStudy (anonymous):

Just use the pythagorean theorem but do 30^2-14^2=a^2 900-196=704 √704=26.53 the length of the ramp along the ground is 26.5 inches along the ground :)

OpenStudy (anonymous):

or it would be 27 inches rounded to the nearest inch :)

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

You're welcome ^.^